# Institutionalizing Folk Theories of Algorithms: How Multi-Channel Networks (MCNs) Govern Algorithmic Labor in Chinese Live-Streaming Industry

### v1

@@ 5.2. Embedding Algorithmic Beliefs into Material and Contractual Structures
This behavior logic aligns precisely with MCNs’ internal understanding of algorithmic unpredictability. Internally, MCNs acknowledge that algorithmic success cannot be reliably manufactured at the individual level. As such, they prefer to spread risk across a broad base of streamers, investing minimally while maximizing returns through volume and service monetization. The more streamers believe in the narrative that effort yields exposure, the more they will spend, regardless of outcome. In this way, the externally constructed folk theory (that visibility is earned through discipline and polish) serves the internal folk theory (that success is probabilistic and unpredictable) by ensuring that MCNs profit from both success and failure, so long as belief is sustained.
It is also important to note that MCNs’ business model depends heavily on a continuous influx of new streamers who are willing to believe in the algorithmic theories MCNs constructs, as well as on retaining those who persist in pursuit of algorithmic success. For MCNs, profitability is sustained not by the success of every individual streamer, but by the broader system’s ability to motivate belief and continued investment. As such, streamers who reject these curated and institutionalized algorithmic theories or choose to exit the system early are largely inconsequential to MCNs’ operational logic. What matters is not universal buy-in, but a sufficient flow of believers willing to participate in the speculative game of algorithmic visibility.
@@ 6. Discussion
This study reveals how folk theories of algorithms—informal beliefs about how algorithmic systems work—are not merely grassroots knowledge among digital workers, but are actively constructed, formalized, and exploited by intermediary institutions. Drawing on nine months of ethnographic fieldwork in China’s live-streaming industry, we show how Multi-Channel Networks (MCNs) strategically develop dual theories about algorithms: internally, they treat algorithms as opaque and unpredictable; externally, they frame algorithms as learnable and responsive. These theories do not simply interpret the algorithm—they organize labor around it. These dual theories of algorithms, what we call internal and external folk theories, form the epistemic foundation upon which MCNs build a regime of algorithmic labor control.
This perspective complements prior live-streaming research in a critical approach. Prior studies have highlighted how live-streaming offers new forms of economic participation, cultural visibility, and affective connection, particularly for marginalized groups such as rural youth, migrant workers, and women (Freeman and Wohn, 2020; Tang et al., 2022; Momin, 2023). MCNs, in turn, have been credited with providing technical support, professional development, and algorithmic literacy, helping new entrants navigate complex platform infrastructures (Zhang and Tong, 2024; Liu et al., 2023; Zhang et al., 2024; Xiao et al., 2024a). We do not dispute these contributions. Indeed, many of our participants expressed appreciation for the opportunities afforded by both platforms and MCNs. For streamers with limited educational backgrounds, rural origins, or unstable offline employment histories, live-streaming was described as one of the few accessible avenues to economic independence and social recognition (Xiao et al., 2024a). MCNs were often credited with lowering entry barriers providing training, equipment, and algorithmic guidance that participants felt they could not have accessed on their own. Yet, our study foregrounds the structural dynamics beneath this perceived empowerment: how organizationally curated algorithmic knowledge disciplines labor without assuming direct managerial responsibility.
@@
Our findings suggest that algorithmic folk theories function not merely as interpretive tools, but as organizational infrastructures through which labor is recruited, directed, and managed. While much of the CSCW literature has emphasized folk theories as emergent, user-driven responses to algorithmic opacity  (Shen et al., 2021; Li et al., 2023; Deng et al., 2023a), our study shifts the analytic focus to institutional actors and their role in constructing and instrumentalizing such theories. To develop this argument, the discussion proceeds in three parts. First, we revisit the concept of folk theories of algorithms, showing how they move beyond grassroots meaning-making to become institutionally manufactured knowledge regimes. We position this as a shift from individual or community-based interpretation to organizational construction (subsection 6.1). Second, we analyze how these institutionally constructed theories serve as instruments of labor governance, enabling MCNs to manage precarity, displace responsibility, and stabilize labor output through soft control (subsection 6.2). Third, we reflect on the broader design and policy implications of these findings. We call for systems and policies that support epistemic transparency, making visible not just what algorithms do, but how knowledge about them is constructed (subsection 6.3).
6.1. From Individual / Community Interpretation of Algorithms to Institutional Construction
A key contribution of this study is to reposition folk theories of algorithms not simply as informal beliefs formed through individual experience, but as forms of knowledge that can be constructed, circulated, and operationalized by organizations for specific managerial and strategic ends. Much of the existing scholarship on folk theories of algorithms has conceptualized them as informal, bottom-up interpretive frameworks that users or communities develop to make sense of opaque algorithmic systems (Bucher, 2018; DeVito et al., 2017; Eslami et al., 2016; Xiao et al., 2024b). These theories are often described as experiential, situated, and emergent, rooted in users’ lived interactions with platforms and shaped through trial and error, peer discussion, and rumor. In this framing, folk theories appear as adaptive tools, helping users navigate digital environments despite uncertainty and asymmetry in platform knowledge. Our study challenges this grassroots-centric view by demonstrating how folk algorithmic knowledge can also be systematically constructed, curated, and institutionalized by organizations. Rather than emerging only from individual experience, folk theories in the context of MCNs are strategically produced and maintained as part of the organization’s epistemic infrastructure. They are stabilized through onboarding routines, coaching scripts, and training materials; circulated via internal handbooks and chats; and selectively modified in response to shifts in algorithmic outcomes. In this process, the algorithm is rendered legible not through spontaneous interpretation, but through deliberate institutional algorithmic theory.
This shift from individuals or communities interpreting algorithmic behavior on their own, to institutions like MCNs actively constructing and distributing standardized interpretations carries several important theoretical implications for CSCW. First, we argue that folk algorithmic knowledge should be reconceptualized as infrastructural, rather than incidental. Previous literature often treats folk theories as emergent, ad-hoc, and peripheral, arising spontaneously when users try to make sense of black-boxed systems. However, our findings show that intermediary organizations such as MCNs deliberately assemble, stabilize, and disseminate particular versions of these theories. They do so not as casual helpers, but as epistemic intermediaries—institutions that claim the authority to explain what the algorithm does, how it works, and what streamers need to do to succeed. In this sense, folk theories of algorithms are not just interpretive stories, but part of the working infrastructure (Pipek and Wulf, 2009; Star and Ruhleder, 1996; Karasti and Blomberg, 2018; Ludwig et al., 2018) that governs digital labor.
@@ 6.1. From Individual / Community Interpretation of Algorithms to Institutional Construction
Second, our findings challenge the prevailing binary in CSCW and HCI that separates algorithmic knowledge into formal (e.g., platform documentation, API guides) and informal (e.g., user speculation, peer discussion, online rumor) categories (Kusk and Nouwens, 2022; Xiao et al., 2024b; Bishop, 2019). In the context of MCNs, this division collapses. The algorithmic knowledge that circulates within these organizations is neither officially sanctioned nor casually improvised. Instead, we conceptualize it as institutionalized folk theory—informal knowledge that is produced and stabilized by intermediary actors who lack direct access to platform-engineered formal knowledge. MCNs are not platform insiders; they operate without formal connections to algorithm developers or privileged access to system logics. Yet they construct and disseminate their own working theories of “how the algorithm works” and embed these into onboarding scripts, coaching sessions, ranking systems, and contractual expectations. These folk theories gain a quasi-formal status—not through technical verification, but through routinized organizational practice, repetition, and discursive reinforcement. In doing so, MCNs transform folk sensemaking into a mechanism of labor governance.
Third, our findings broaden CSCW’s understanding of how algorithmic knowledge circulates in everyday life. Prior work has shown that workers and users build collective knowledge through grassroots means like online forums, creator communities, or worker-led organizing (DeVito et al., 2017; Karizat et al., 2021; Xiao et al., 2024b). These are important, but they are not the only sites where algorithmic knowledge is produced. Our study shows that intermediary institutions like MCNs also serve as powerful knowledge brokers. They don’t just transmit information but also shape and curate it. They decide which explanations of the algorithm get repeated, which practices get standardized, and which beliefs get embedded into training and labor management. Unlike peer-to-peer communities, MCNs combine technical interpretation with organizational authority. They have the resources and legitimacy to institutionalize speculative beliefs into widely accepted norms. In doing so, they turn uncertainty into order, and speculation into action, making algorithmic systems workable, even if not fully knowable.
Our study thus pushes CSCW to look beyond individuals or user groups when studying how people interact with algorithmic systems. It invites scholars to future consider the organizational infrastructures that mediate algorithmic knowledge and how those infrastructures shape not only what workers believe, but how they work, comply, and sustain hope.
6.2. Folk Theories of Algorithms for Labor Management

### v2

@@ 5.2. Embedding Algorithmic Beliefs into Material and Contractual Structures
This behavior logic aligns with MCNs’ internal understanding of algorithmic unpredictability. Internally, MCNs acknowledge that algorithmic success cannot be reliably manufactured at the individual level. As such, they prefer to spread risk across a broad base of streamers, investing minimally while maximizing returns through volume and service monetization. The more streamers believe in the narrative that effort yields exposure, the more they will spend. In this way, the externally constructed algorithmic interpretations (that visibility is earned through discipline and polish) serves the internal interpretations (that success is probabilistic and unpredictable) by ensuring that MCNs profit from both success and failure, so long as belief is sustained.
6. Reactions to MCNs’ Discursive Construction of Algorithms: Live-Streamers’ Compliance, Doubt, and Departure (RQ3)
In the preceding sections, we examined how MCNs construct algorithmic interpretations (section 4) and how these interpretations are mobilized as mechanisms of labor control (section 5). It is important to clarify that MCNs’ capacity to govern live-streamers’ labor rests largely on algorithmic discourse and, crucially, on streamers’ acceptance of that discourse. On most major Chinese platforms, live-streamers are technically able to broadcast independently without affiliating with MCNs. Live-streaming itself requires only minimal equipment, and there is no inherent necessity to invest in professional devices or studio infrastructures. In practice, however, many streamers enter the live-streaming industry after encountering MCNs’ promotional narratives with optimistic career expectations. After joining, these expectations are continuously reinforced through everyday guidance and coaching. Carried by this promise of future success, large numbers of streamers enter the industry, collaborate with MCNs, continue to invest their labor, and willingly upgrade their equipment in pursuit of better streaming performance and increased algorithmic visibility.
Because labor control is mediated through belief rather than coercion, streamers’ responses to MCNs’ algorithmic interpretations are not uniform. Some comply with these expectations and internalize them as professional norms (subsection 6.1); others begin to question their validity as promised outcomes fail to materialize (subsection 6.2); still others ultimately exit both MCNs and the live-streaming industry altogether (subsection 6.3). In what follows, we analyze these divergent responses to show how algorithmic discourse is variably taken up, negotiated, and resisted in streamers’ everyday labor practices.
6.1. Compliance with MCNs’ Labor Management
Compliance is the most common response among live-streamers in our study. For many participants, accepting MCNs’ algorithmic interpretations represents the least psychologically and cognitively demanding option under conditions of algorithmic uncertainty. Rather than constantly questioning platform mechanisms or organizational guidance, compliance offers a sense of direction and emotional relief in an otherwise opaque and volatile environment.
This tendency toward compliance is closely tied to the demographic composition of China’s live-streaming industry. The industry is heavily female-dominated and disproportionately attracts young women from rural areas, many of whom face constrained employment opportunities in other sectors of the labor market (Xiao et al., 2024a). For these workers, live-streaming often appears to be one of the few accessible pathways to flexible income generation or upward mobility. Within this context, MCNs’ narratives of algorithmic fairness, effort-based success, and eventual recognition become especially compelling.
For many streamers, compliance is therefore not a passive or naïve choice, but a pragmatic response to limited alternatives. As several participants emphasized, when everyday life is marked by economic precarity and uncertainty, any option that promises even a modest chance of improvement can feel worth pursuing. This pattern has been documented in prior research on China’s live-streaming industry (Xiao et al., 2024a) and was also evident in our ethnographic fieldwork. Multiple participants (e.g., L1, L4, L6, L13, L16) noted that even if they left live-streaming, they had no clear sense of what other work they could realistically do.
This sense of constrained choice is also reflected in the demographic profiles of our participants (Table 1), particularly in relation to their educational backgrounds and earnings. L6, who completed only high school, described her perceived options as extremely limited: “Working in an electronics factory is exhausting, but doing food delivery or driving for Didi requires renting a car, which is very expensive up front—and besides, both industries are dominated by men.” In such circumstances, complying with MCNs’ guidance and continuing to invest effort in live-streaming can appear more reasonable than exiting the industry altogether. [^1] It is important to note that these findings do not contradict prior HCI research highlighting the empowering potential of live-streaming for marginalized groups (Tang et al., 2022; Wang et al., 2024). Many participants in our study also described tangible improvements in their living conditions after becoming streamers. For instance, L6 explained that she was able to move from a crowded dormitory to a small studio apartment. Although both housing situations were modest by urban standards, this change still represented a meaningful improvement in her daily life. In this sense, exploitation and improvement coexist: live-streaming may expose workers to new forms of control while simultaneously offering incremental material benefits. Rather than opposing earlier accounts, our findings show how these dynamics are experienced together in streamers’ everyday lives. For some participants, the fact that live-streaming allowed them to live slightly better than before also made its exploitative aspects feel less severe. As L6 put it, “other jobs might be even worse.”
6.2. Doubting MCNs’ Algorithmic Narratives
Doubt emerges when live-streamers begin to notice persistent discrepancies between MCNs’ algorithmic explanations and their own lived outcomes. As promised visibility fails to materialize despite sustained effort and investment, some streamers start to question whether the algorithm is as effort-sensitive as they were led to believe.
These doubts are often triggered by repeated experiences of stagnation or regression. Streamers described situations in which they followed MCNs’ guidance closely—maintaining long streaming hours, upgrading equipment, adjusting appearance, and optimizing interaction—yet observed little change in traffic or recommendations. Over time, such experiences erode confidence in the causal links promoted by MCNs’ narratives. As L9 put it, “I did everything they told me to do. I streamed longer, bought better lights, smiled more. But the numbers didn’t move. After a while, I started wondering if it really worked the way they said.”
Doubt does not necessarily lead to immediate resistance or exit. For many streamers, questioning algorithmic explanations coexists with continued compliance. Streamers may privately suspect that MCNs’ guidance is overstated or selectively framed, while still following it in practice. This ambivalence reflects the high stakes of withdrawal: exiting the industry means giving up not only future income but also the possibility that success might still arrive. As L12 explained, “I knew something was off, but I kept going. If I stopped, then all the effort before would really be wasted.”
Doubt is also shaped by asymmetric access to information. Streamers rarely have alternative explanations for poor performance beyond those provided by MCNs. Platform algorithms remain opaque, and peer comparison often reinforces uncertainty rather than clarity. When other streamers occasionally succeed, their success is frequently interpreted as confirmation that the system still works, rather than as evidence of randomness. This keeps doubt fragmented and individualized, rather than developing into collective critique.
As a result, doubt functions less as a direct challenge to MCNs’ authority and more as a fragile, unstable state. Streamers oscillate between belief and skepticism, hope and frustration, without a clear pathway toward resolution. This prolonged state of uncertainty allows MCNs’ algorithmic discourse to continue exerting influence, even as its credibility weakens. Doubt thus represents an intermediate response to algorithmic control—one that signals cracks in belief while still binding streamers to continued participation.
6.3. Departure from the Live-Streaming Industry
Departure represents the point at which belief in MCNs’ algorithmic interpretations becomes untenable and continued participation can no longer be justified. Unlike compliance or doubt, departure involves a concrete break from both MCNs and, in many cases, the live-streaming industry itself. This decision typically follows prolonged periods of stagnation, accumulated financial pressure, and emotional exhaustion.
Notably, departure is rarely immediate. Several participants described an intermediate phase in which they disengaged from MCNs while continuing to stream independently. F1 and F7 both left their MCNs and attempted to live-stream on their own for a period before ultimately exiting the industry altogether. These transitions were motivated by a desire to test whether success was possible outside MCNs’ control and guidance. As F1 explained, “I wanted to see if it was really the MCN holding me back, or if I just needed more time. So I kept streaming on my own.” When performance did not improve despite this shift, the promise of algorithmic fairness in live-streaming lost its credibility.
At the same time, independent streaming does not necessarily preclude success. F7, for example, accumulated approximately 28,000 followers after leaving his MCN and streaming independently, with most of this audience growth occurring outside the MCN system. His case illustrates that algorithmic visibility is not inherently dependent on MCN affiliation. A small number of independent streamers are able to develop their own understandings of platform dynamics and cultivate relatively stable audiences without organizational mediation.
However, such cases are exceptional rather than typical in China, given the generally low level of digital literacy among most Chinese streamers. Independent success requires not only time and sustained experimentation, but also the capacity to develop one’s own folk theories about algorithmic visibility, audience engagement, and platform rhythms. While independent success remains possible in principle, it remains far from accessible to most streamers in practice. As M1 explained, this disparity also helps account for the prominence of MCNs in China: “MCNs may be less prevalent in Western contexts because most streamers there have relatively high digital literacy, but that is not the case in China.”
Other streamers attempted to resolve doubt by switching organizational affiliations. F5 moved between two different MCNs before leaving the industry, hoping that a change in resources or guidance might lead to different outcomes. This strategy reflects continued belief in the underlying algorithmic narrative: that success remains achievable under the right conditions. Only after repeated disappointment across organizational contexts did exit become a viable option.
These trajectories suggest that departure is not a rejection of live-streaming from the outset, but the outcome of sustained efforts to make the system work. Streamers often exhaust available strategies—compliance, optimization, organizational switching, and independent streaming—before concluding that further investment is unlikely to pay off. By the time they leave, many describe a sense of clarity mixed with resignation. As F7 put it, “At some point, I realized it wasn’t about working harder or finding the right MCN. I just couldn’t keep going like this.”
We intentionally sampled streamers with different exit trajectories in order to capture this processual nature of departure. Departure marks the limit of MCNs’ discursive control: when hope, optimization, and even organizational switching no longer sustain participation, streamers disengage from the industry altogether.
To conclude, compliance, doubt, and departure reveal a contextual condition shaping MCNs’ discursive influence. Many streamers in our study have limited formal education and digital literacy, constraining their ability to independently interpret platform algorithms. Under such conditions, MCNs’ algorithmic explanations become the primary resources for making sense of uncertainty and performance outcomes. This constraint produces a form of constrained agency. Streamers voluntarily rely on MCNs’ narratives as a pragmatic strategy for navigating opaque platforms, while limited alternatives make such reliance difficult to avoid. Trust is sustained not through overt coercion, but through asymmetric access to algorithmic knowledge and interpretive authority. As long as MCNs’ explanations remain credible, they organize labor by shaping how effort, investment, and future success are understood. When credibility erodes, streamers move from compliance to doubt and ultimately to departure, marking the withdrawal of belief that underpins MCN-managed labor.
@@ 7. Discussion
This study reveals how discursive construction of algorithms—informal beliefs about how algorithmic systems work—are actively constructed, formalized, and exploited by intermediary institutions. Drawing on nine months of ethnographic fieldwork in China’s live-streaming industry, we show how Multi-Channel Networks (MCNs) strategically develop dual interpretations about algorithms: internally, they treat algorithms as opaque and unpredictable; externally, they frame algorithms as learnable and responsive. These interpretations do not simply interpret the algorithm—they organize labor around it.
Our ethnographic findings illuminate the organizational dynamics that underpin these opportunities. Many MCNs operate under sustained pressure to recruit new streamers and to keep existing streamers active. In practice, MCNs often take a proactive stance toward streamers, investing considerable effort in outreach, onboarding, and persuasion. Streamers described signing with MCNs as a relatively flexible and low-commitment decision: many felt able to choose among agencies they preferred, and exit costs were generally low, both contractually and financially. Continued streamer participation nonetheless remains a central condition for MCNs’ own survival and profitability. This organizational dependence shapes how MCNs engage with streamers over time. Rather than relying primarily on coercive control, MCNs engage in ongoing discursive work to sustain motivation and align streamer expectations with organizational needs. Algorithmic interpretations play a central role in this process.
Concretely, MCNs circulate an external-facing algorithmic narrative built around two claims: (1) the algorithm is meritocratic and rewards persistence; (2) visibility follows identifiable signals that can be optimized; (RQ1). They translate the promised signals into mandatory-like routines (fixed schedules, performance targets, appearance standards, and continuous affective display) and then tie compliance to material arrangements that generate revenue for MCNs. Profit comes from commissions on successful streamers’ earnings, plus more reliable income from services sold to a large pool of newcomers, including training packages, equipment rental or sales, studio fees, and platform-operation support. The same narrative that attributes success to optimization also legitimizes recurring purchases and escalating investment (RQ2). Streamers’ responses cluster into three trajectories. Many comply because the narrative offers direction and reduces cognitive burden under opacity; compliance also feels low-risk given flexible contracts and low exit penalties. Others express doubt when repeated effort and spending fail to change traffic, yet doubt often coexists with continued compliance due to sunk costs and the hope of a turnaround. Departure occurs when low returns, debt, and exhaustion accumulate to the point where the narrative no longer sustains continued participation, sometimes after a short phase of independent streaming to test whether MCN mediation was the bottleneck (RQ3).
@@
Our findings suggest that MCNs’ discursive construction of algorithms functions not merely as interpretive framework, but as organizational infrastructures through which labor is managed. While much of the CSCW literature has emphasized algorithmic interpretation as emergent, user-driven responses to algorithmic opacity (e.g., folk theories)  (Shen et al., 2021; Li et al., 2023; Deng et al., 2023), our study shifts the analytic focus to institutional actors and their role in intentionally constructing and instrumentalizing such algorithmic interpretations. To develop this argument, the discussion proceeds in three parts. First, we pay attention to the organizational discursive construction of algorithms in the platform economy (subsection 7.1). Second, we analyze how these discursive construction of algorithms serve as instruments of labor governance, enabling MCNs to manage precarity, displace responsibility, and stabilize labor output through soft control (subsection 7.2). Third, we reflect on the broader design and policy implications of these findings. We call for systems and policies that support epistemic transparency, making visible not just what algorithms do, but how knowledge about them is constructed (subsection 7.3).
7.1. Constructing Algorithmic Authority: MCNs’ Discursive Construction of Algorithms
A key contribution of this study is to shift attention from user-centered interpretations of algorithms to the organizational production of algorithmic meaning. Prior scholarship on algorithmic interpretations, including work on folk theories, has largely examined how users or communities develop informal, bottom-up understandings to make sense of opaque algorithmic systems (Bucher, 2018; DeVito et al., 2017; Eslami et al., 2016; Xiao et al., 2024b). These interpretations are typically described as experiential and emergent, shaped through everyday interaction, peer exchange, and speculation. Within this framing, algorithmic interpretations function as practical sensemaking resources that help users cope with uncertainty and informational asymmetries.
In contrast, our study focuses on how algorithmic interpretations are discursively constructed by organizations. In the context of MCNs, algorithmic knowledge does not primarily arise from individual trial and error, but is strategically produced, stabilized, and circulated through organizational discourse. These interpretations are embedded in routine training, guidance, and evaluation practices, forming part of the epistemic infrastructure through which MCNs manage and coordinate streamer labor. This shift in analytic focus highlights that algorithmic knowledge is not merely incidental or interpretive, but infrastructural. MCNs act as epistemic intermediaries that claim authority over how algorithms should be understood and acted upon, translating platform opacity into actionable narratives aligned with organizational goals. As such, the discursive construction of algorithms operates as a form of working infrastructure (Pipek and Wulf, 2009; Star and Ruhleder, 1996; Karasti and Blomberg, 2018; Ludwig et al., 2018) that shapes labor practices, expectations, and investment, rather than simply reflecting users’ interpretations.
Importantly, the epistemic authority of MCNs does not come from possessing “correct” or privileged knowledge about platform algorithms. Their understandings are partial, heuristic, and contingent. MCN staff themselves routinely acknowledged that their recommendations may work in some cases yet fail in others, and that they cannot predict outcomes for any specific streamer with confidence. MCNs’ power instead comes from their ability to author and institutionalize algorithmic interpretations under conditions of opacity. In this sense, labor control is generated less by “knowing the algorithm” and more by controlling the narratives through which algorithmic outcomes are explained, attributed, and acted upon. Without the infrastructuralization of algorithmic explanations, MCNs would struggle to sustain their organizational operations, lacking a credible basis for recruitment, retention, and revenue generation, rendering their intermediary role difficult to maintain. While existing scholarship on folk theories has richly documented grassroots sensemaking under algorithmic opacity, it has rarely examined how algorithmic explanations become institutionalized as durable cognitive infrastructures. Our findings foreground this organizational transformation of algorithmic interpretation and its consequences for labor governance.
@@ 7.1. Constructing Algorithmic Authority: MCNs’ Discursive Construction of Algorithms
Second, our findings challenge the prevailing binary in CSCW and HCI that separates algorithmic knowledge into formal (e.g., platform documentation, API guides) and informal (e.g., user speculation, peer discussion, online rumor) categories (Kusk and Nouwens, 2022; Xiao et al., 2024b; Bishop, 2019). In the context of MCNs, this division collapses. The algorithmic knowledge that circulates within these organizations is neither officially sanctioned nor casually improvised. Instead, we conceptualize it as MCNs’ discursive construction of algorithms—informal knowledge that is produced and stabilized by intermediary actors who lack direct access to platform-engineered formal knowledge. MCNs are not platform insiders; they operate without formal connections to algorithm developers or privileged access to system logics. Yet they construct and disseminate their own working interpretations of “how the algorithm works” and embed these into onboarding scripts, coaching sessions, ranking systems, and contractual expectations. These discursive constructions gain a quasi-formal status—not through technical verification, but through routinized organizational practice, repetition, and discursive reinforcement. In doing so, MCNs transform folk interpretation into a mechanism of labor governance.
Third, a key feature of these discursive constructions lies in the tension between MCNs’ organizational practices and their public claims. As companies, MCNs operate under conditions of limited control over algorithmic outcomes. Internally, they manage risk through batch recruitment, uneven resource allocation, and service-based monetization that does not depend on any individual streamer’s success. Externally, however, MCNs present themselves as capable algorithmic guides. This contrast allows MCNs to reconcile uncertainty with authority: organizational practices are oriented toward probabilistic returns and volume-based profit, while external discursive claims foreground individualized improvement and attainable success for steamers. The gap between what MCNs do as firms and what they claim about the algorithm is not incidental. It is precisely this gap that enables discursive constructions to function as labor-governing infrastructure, sustaining participation and legitimizing continued investment despite unstable outcomes.
Moreover, we argued that the belief in MCNs’ algorithmic interpretations reflects unequal capacities for knowledge production within platformized labor markets. In contrast to MCNs’ discursive construction of algorithms, prior research shows that experienced content creators can develop their own informal understandings of platform behavior through prolonged engagement and experimentation (DeVito, 2022; Wu et al., 2019; Moran et al., 2022). Our ethnographic data indicate that most streamers in MCN-managed live-streaming ecosystems lack the material and epistemic conditions required to develop comparable independent interpretations. Many participants entered live-streaming with limited prior experience in content production, unstable offline employment, and strong pressure to generate income quickly. Under these conditions, systematic experimentation with content strategies or algorithmic responses was difficult to sustain. MCNs occupy this gap by offering ready-made explanations of how the algorithm works and how effort should be organized. These explanations are persuasive not because they are demonstrably accurate, but because they provide a coherent account that reduces uncertainty and distributes responsibility. For streamers who lack the resources to independently test or verify algorithmic behavior, trusting MCNs’ interpretations becomes a rational response to structural constraints rather than a simple failure of critical judgment.
7.2. Discursive Construction of Algorithms for Labor Management